# No more hard prompts: SoftSRV prompting for synthetic data generation

### v2

@@ Appendix B Hard Prompt Templates
Below, we report the templates used for hard-prompting baselines for each benchmark dataset. Figure 6 provides the question templates while Figure 7 shows the answer templates. Figure 8 reports the critique and refine templates for $\mathrm{HP}_{\mathrm{SR}}$. Figure 9 provides the template for the undiversified $\mathrm{HP}$ method, described in Appendix A.5, for the GSM8K benchmark .
MBPP Question Template:
@@
Consider the following python question: [insert example question] Now generate 10 different questions that require writing a Python function similar to the example above. Make sure each question is different and sufficiently rephrased. Please make sure you generate questions, and not answers. Please make sure each question you generate has a well-defined answer. Question 1:
@@
Figure 6: Question template for MBPP, GSM8K and BoolQ benchmarks for the $\mathrm{HP}$ method.
@@
Figure 7: Answer template for MBPP, GSM8K and BoolQ benchmarks for the $\mathrm{HP}$ and $\mathrm{HP}_{\mathrm{SR}}$ method.
@@
Figure 8: Refine template for MBPP, GSM8K and BoolQ benchmarks for the $\mathrm{HP}_{\mathrm{SR}}$ method.
GSM8K Question Template for undiversified $\mathrm{HP}$: Please generate a question that requires solving a grade-school math problem. Here is an example of such a question: [insert example question] Now generate a new question. Please make sure your question is not too similar to the example above. Please make sure you generate a question, and not an answer. Please make sure the question you generate has a well-defined answer.
Figure 9: Question template for the undiversified $\mathrm{HP}$ method for the GSM8K benchmark.

### v3

@@ Appendix C Natural Language Prompt Templates
Below, we report the templates used for prompt template baselines for each benchmark dataset. Figure 8 provides the question templates while Figure 9 shows the answer templates. Figure 10 provides the template for the undiversified $\mathrm{PT}$ method, described in Appendix A.4, for the GSM8K benchmark. Figure 11 reports the critique and refine templates for $\mathrm{PT}_{\mathrm{SR}}$.
Appendix D Generated Text Examples
In Figure 12, we provide examples of synthetic questions for GSM8K (math questions), MBPP (Python questions) and BoolQ (passage and question pairs) generated using the $\mathrm{SS}_{\mathrm{MC}}$ method. Notice that the questions generated for BoolQ require longer range dependencies and may be more complex to generate and answer. Finally, for comparison, in Figure 13, we show examples from the train split of each dataset. These train set questions appear similar in style to the questions generated by the $\mathrm{SS}_{\mathrm{MC}}$ method.
@@ Appendix D Generated Text Examples
MBPP Question Template: Consider the following python question: [insert example question] Now generate 10 different questions that require writing a Python function similar to the example above. Make sure each question is different and sufficiently rephrased. Please make sure you generate questions, and not answers. Please make sure each question you generate has a well-defined answer. Question 1:
@@
Figure 8: Question template for MBPP, GSM8K and BoolQ benchmarks for the $\mathrm{PT}$ method.
@@
Figure 9: Answer template for MBPP, GSM8K and BoolQ benchmarks for the $\mathrm{PT}$ and $\mathrm{PT}_{\mathrm{SR}}$ method.
GSM8K Question Template for undiversified $\mathrm{PT}$: Please generate a question that requires solving a grade-school math problem. Here is an example of such a question: [insert example question] Now generate a new question. Please make sure your question is not too similar to the example above. Please make sure you generate a question, and not an answer. Please make sure the question you generate has a well-defined answer.
Figure 10: Question template for the undiversified $\mathrm{PT}$ method for the GSM8K benchmark.
@@
Figure 11: Refine template for MBPP, GSM8K and BoolQ benchmarks for the $\mathrm{PT}_{\mathrm{SR}}$ method.
GSM8K Generated Text Examples
     Julian is twice as old as his little brother, Eli. In  12 years, Julian will be 37. How old is Eli today?
     Sarah used 1/3 of her grass cutting saw’s fuel in  cutting the grass in front of her house and used 2/5  of the remaining fuel to cut the grass in her back  yard. If she started with 30 liters of fuel, how much  fuel did she use in cutting the front yard?
     A boy is jogging at 4 m/s. How long will he take to  cover a distance of 5 km?
MBPP Generated Text Examples
     Write a python function to find the average of three  given numbers.
     Given a sequence of characters, create a new sequence  such that every character is repeated k times.
     Write a function that takes in two numbers and returns  the sum of the two numbers.
BoolQ Generated Text Examples
     PASSAGE: There are no known cases of anyone catching  a common cold by exposure to the cold weather. The  most likely cause of a common cold is a virus. The  symptoms of the common cold can include a runny nose,  sore throat, cough, sneezing, nasal congestion, and  watery eyes. A common cold is considered to be mild  when only mild symptoms are present. Severe symptoms  include sinus pain and fever.  QUESTION: can the common cold be caused by cold  weather
     PASSAGE: Land Before Time V: The Mysterious Island  (1997) is an American animated direct-to-video film.  It is the fifth film in The Land Before Time series  and the first film in the series to be animated in  Canada and also the first film in the series to be  created by Universal Cartoon Studios.  QUESTION: is the land before time 5 a sequel
     PASSAGE: The semilunar valves of the heart (or  sometimes called the sigmoid valves) are the valves  that prevent backflow from the aorta and pulmonary  artery. They are situated at the beginning of these  arteries. The pulmonary semilunar valve is on the  right side of the heart, while the aortic semilunar  valve is on the left. They both have three leaflets,  each of which is attached to the arterial wall by the  corpora fibrosa. They are attached to the artery in  such a manner that they cannot fold over to block the  flow of blood.  QUESTION: are the aortic and pulmonary valves  semilunar valves
Figure 12: Examples of generated synthetic questions generated by $\mathrm{SS}_{\mathrm{MC}}$.
GSM8K Example Questions From Train Set
     A choir was singing a song that involved 30 singers.  In the first verse, only half of them sang. In the  second verse, a third of the remaining singers joined  in. How many people joined in the final third verse  that the whole choir sang together?
     Kyle is 5 years older than Julian. Julian is 20 years  younger than Frederick. Frederick is 2 times older  than Tyson. If Tyson is 20, how old is Kyle?
     Mr. Williams bought 10 gallons of juice for a party.  Each gallon has 10 cups. At the party, 5 cups of juice  were left. How many cups of juice were drunk?
MBPP Example Questions From Train Set
     Write a function to move all the numbers in it to the  given string.
     Write a function to find the largest subset where each  pair is divisible.
     Write a function to increment the numeric values in  the given strings by k.
BoolQ Example Questions From Train Set
     PASSAGE: Since being fixed on the fourth Thursday in  November by law in 1941, the holiday in the United  States can occur on any date from November 22 to 28.  When it falls on November 22 or 23, it is not the  last Thursday, but the penultimate Thursday in  November. Regardless, it is the Thursday preceding  the last Saturday of November.  QUESTION: is thanksgiving the last thursday of  november every year
     PASSAGE: The Rocky Mountains, also known as the  Rockies, are a major mountain range in western North  America. The Rocky Mountains stretch more than 3,000  miles (4,800 km) from the northernmost part of British  Columbia, in western Canada, to New Mexico, in the  Southwestern United States. Within the North American  Cordillera, the Rockies are somewhat distinct from the  Pacific Coast Ranges, Cascade Range, and the Sierra  Nevada, which all lie further to the west.  QUESTION: is the sierra nevada part of the rocky  mountains
     PASSAGE: The first player to get rid of their last  card (‘‘going out’’) wins the hand and scores points  for the cards held by the other players. Number cards  count their face value, all action cards count 20, and  Wild and Wild Draw Four cards count 50. If a Draw Two  or Wild Draw Four card is played to go out, the next  player in sequence must draw the appropriate number of  cards before the score is tallied.  QUESTION: can you finish a uno game on a wild card
Figure 13: Examples of questions from the train sets.